Find the number of combinaTIONS? 7c5
Answer by stanbon(75887)

Find the number of combinaTIONS? 7c5
7C5 = 7!/[(7-5)!*5!]
= (7*6)/(1*2)
= 7*3 = 21
================
Cheers,
Stan H.